the width of a triangle i 3 inches less than its length, and the area is 108 square inches.what are the length and width of the rectangle?
Answer by stanbon(75887)

the width of a triangle is 3 inches less than its length, and the area is 108 square inches.what are the length and width of the rectangle?
--------
Let length be L
Then width = L-3
----
Area = length*width
108 = L(L-3)
L^2 - 3L - 108 = 0
(L-12)(x+9) = 0
Positive solution:
Length = 12 inches
Width = L-3 = 9 inches
---------------------------
Cheers,
Stan H.
================